# Diffusion-based Dynamic Contract for Federated AI Agent Construction in Mobile Metaverses

### v1

@@ V-C Enhanced Diffusion Policy for Dynamic Contract Design
Figure 7: Performance of the EDMSAC algorithm in optimal dynamic contract design under different parameter configurations.
VI Simulation Results
VI-A Experimental Setup
Experiment Settings: We consider ESs with homogeneous hardware capabilities and classify them into two types: high WTP and low WTP [19, 29]. Specifically, $\theta^{t}_{1}$ and $\theta^{t}_{2}$ are randomly sampled within $[15,20]$ and $[20,25]$, respectively. $p^{1}_{k}$ and $p^{2}_{k}$ are randomly generated following the Dirichlet distribution [4]. The setting of simulation parameters are summarized in Table I [33, 32, 15, 16, 22]. For the algorithm design, DM-based actor networks use the Variance Proportional (VP) noise schedule strategy, and the actor and critic networks are trained by the Adam optimizer, with a weight decay rate of $0.0001$ for regularizing model weights [36]. In addition, the discount factor $\gamma$ is set to $0.99$, and $\tau$ is set to $0.005$ to soft update the target networks. Notably, our experiments use PyTorch with CUDA 12.0 on an NVIDIA GeForce RTX 3080 Laptop GPU.
Scheme Details: To evaluate the performance of the proposed scheme, we compare it with other typical schemes:
VI-B Performance Evaluation of Mobile AI Agents
As illustrated in Fig. 4, we present two case studies in which a mobile AI agent operates a real smartphone to perform tasks. For the perception module, we leverage Qwen-VL-Plus to create it, and we use the ConvNextViT-document model and the GroundingDINO model as text and icon recognition tools, respectively[30]. For the decision-making, reflection, and planning modules, we leverage Qwen-VL-Max to create them. To evaluate the performance of the constructed AI agent, we adopt several evaluation metrics: Decision Accuracy (DA), which measures the accuracy of the decision-making module; Reflection Accuracy (RA), which assesses the accuracy of the reflection module; and Planning Accuracy (PA), which evaluates the accuracy of the planning module. As shown in Fig. 5, the performance of the AI agent in executing simple tasks, such as Case 1, is superior to that in handling more complex tasks, like Case 2, as evidenced by higher values of DA, RA, and PA. Notably, when executing simple tasks, the AI agent may perform additional rounds to verify task completion, leading to increased service latency.
@@ VI-B Performance Evaluation of Mobile AI Agents
VI-C Performance Evaluation of the EDMSAC Algorithm
In Fig. 6, we present a performance comparison between the proposed EDMSAC algorithm and several representative DRL algorithms in optimal dynamic contract design, including DMSAC [36], DM-based Deep Deterministic Policy Gradient (DMDDPG) [28], PPO, SAC, and DDPG. From Figs. 6(a) and 6(b), we observe that the proposed EDMSAC algorithm outperforms other DRL algorithms in optimal dynamic contract design because DRL algorithms struggle in different suboptimal solution spaces, especially PPO algorithms. The superior performance of the proposed EDMSAC algorithm can be primarily attributed to two key aspects. First, the EDMSAC algorithm has a strong capability to capture complex environmental information because of DMs [28], which helps avoid the identification of suboptimal dynamic contracts. Moreover, its fine-grained policy tuning mitigates the impact of environmental randomness, enhancing decision robustness. Second, through structural pruning techniques, the unimportant neurons in the DM-based actor networks are masked rather than permanently removed, thereby preserving the network structure while reducing computational redundancy. Therefore, the EDMSAC algorithm can achieve higher denoising efficiency, enabling the generation of more optimal and stable dynamic contracts. From Fig. 6(c), we also observe that the EDMSAC algorithm achieves a higher total profit of the cloud server compared with other DRL algorithms. Overall, the above analysis demonstrates the effectiveness and superiority of the proposed EDMSAC algorithm in optimizing dynamic contract design, which enhances the system performance of federated AI agent construction under dynamic information asymmetry.
As shown in Fig. 7, we present the performance of the EDMSAC algorithm in optimal dynamic contract design with different parameter configurations. Fig. 7(a) illustrates the performance of the EDMSAC algorithm under different pruning rates, i.e., $30\%$, $50\%$, and $70\%$. We observe that the EDMSAC algorithm demonstrates superior performance under higher pruning rates, leading to faster convergence and higher test rewards. The reason is that a higher pruning rate indicates that more redundant and unimportant neurons in the DM-based actor networks are structurally masked. As a result, the pruned DM-based actor networks exhibit better generalization capabilities and are more effective at learning essential features in optimal dynamic contract design. Fig. 7(b) presents the performance of the EDMSAC algorithm under different denoising steps, i.e., $3$, $6$, and $12$. We observe that the EDMSAC algorithm achieves faster convergence and yields higher total profits for the cloud server under a smaller denoising step. The reason is that a smaller denoising step can achieve higher computational efficiency of the DM-based actor networks. Fig. 7(c) illustrates the performance of the EDMSAC algorithm under two groups of different learning rates. The first group is the actor networks with a learning rate of $2\times 10^{-7}$ and the critic networks with a learning rate of $2\times 10^{-6}$. The second group is the actor networks and critic networks with the same learning rate of $1\times 10^{-4}$. We observe that the EDMSAC algorithm has better performance under the first group of learning rates, indicating that the pruned actor networks need to learn an efficient policy in optimal dynamic contract design under small learning rates.
VI-D Performance Evaluation of the Dynamic Contract Scheme
In Fig. 8, we illustrate the two-period dynamic contracts generated by the EDMSAC algorithm under three network states, with a pruning rate of $0.5$ and a denoising step of $6$. Additionally, the learning rates of the actor and critic networks are set to $2\times 10^{-7}$ and $2\times 10^{-6}$, respectively. Attributed to the exploration and generalization capabilities of DMs, the proposed EDMSAC algorithm demonstrates the ability to generate optimal dynamic contracts across varying network states, indicating that our algorithm possesses strong adaptability to the dynamic environments inherent in federated AI agent construction, thereby ensuring robust system performance in mobile metaverses. We observe that in period 2, ESs with higher types perform more training rounds $T_{2}^{k}$ for agent module creation and consequently receive higher rewards $R_{2}^{k}$, which validates Lemmas 1 and 2, respectively. In period 1, ESs with higher types also perform more training rounds $T_{i}^{k}$ for agent module creation and receive higher rewards $R_{i}^{k}$, which demonstrates Lemmas 7 and 8, respectively.
@@ VI-D Performance Evaluation of the Dynamic Contract Scheme
In Fig. 9, we compare the proposed two-period dynamic contract scheme with three baseline contract schemes. Fig. 9(a) presents the test reward comparison between the proposed scheme and three baseline contract schemes. Our scheme consistently achieves higher test rewards compared with the other three contract schemes, indicating that the proposed two-period dynamic contract scheme possesses superior performance. In addition, the test rewards obtained by dynamic contract schemes consistently surpass those of static contract schemes, highlighting the advantages of incorporating adaptability in contract design. The reason is that, unlike static contract schemes, dynamic contract schemes take into account the profit of the cloud server in the second period. As a result, the total profit of the cloud server achieved under the dynamic contract scheme is higher than that of the static contract scheme. Fig. 9(b) illustrates the relationship between the total profit of the cloud server and varying numbers of ESs under four contract schemes. No matter under which contract scheme, as the number of ESs increases, the total profit of the cloud server also increases. Moreover, compared with the three baseline contract schemes, the proposed two-period dynamic contract scheme can achieve the highest total profit for the cloud server. This is primarily because the proposed scheme is capable of designing suitable contracts for ESs based on their respective types rather than generating contracts randomly, which enables more effective resource utilization for federated AI agent construction and contributes to higher total profits for the cloud server. Fig. 9(c) presents the impacts of the scaling factor $\alpha$ and the discount factor $\beta$ on the total profit of the cloud server under different numbers of ESs. When $\beta$ remains constant, the total profit of the cloud server increases with the rise of $\alpha$. The reason is that a larger $\alpha$ reflects a greater overall magnitude of economic benefits allocated to the cloud server, thereby enhancing its total profit. In addition, when $\alpha$ remains constant, the total profit of the cloud server increases as $\beta$ increases. This is primarily because the cloud server can obtain higher profits in the second period of federated AI agent construction, which also enhances its total profit.
VII Conclusion
@@
In this paper, we have studied federated AI agent construction in mobile metaverses. Specifically, we have proposed an edge-cloud collaboration-based federated AI agent construction framework in mobile metaverses. In this framework, ESs act as agent infrastructures to create agent modules in a distributed manner, and then the cloud server integrates them into complete AI agents and deploys the constructed AI agents at the edge, thereby providing AI agent services for users in mobile metaverses. Considering the dynamic information asymmetry in federated AI agent construction, we have designed a two-period dynamic contract model to incentivize ESs to continuously participate in federated AI agent construction. Furthermore, to effectively identify optimal dynamic contracts, we have proposed the EDMSAC algorithm, which dynamically prunes redundant neurons in the DM-based actor networks, thereby enhancing denoising efficiency for dynamic contract sampling. Extensive simulations have demonstrated the effectiveness and robustness of the proposed scheme and the EDMSAC algorithm. For future work, we aim to develop a mobile AI agent paradigm built upon the proposed framework, which will empower AI agents with enhanced adaptability and decision-making capabilities in real-world mobile environments.
Appendix A Proof for Theorem 2
@@
We use the contradiction method to demonstrate the uniqueness of the optimal reward in period 2. We assume that there is another optimal reward $\hat{R}_{i}^{2}$ that facilitates the greater profit of the cloud server. Since the utility of the cloud server is inversely proportional to the total reward, we can obtain $\sum_{i=1}^{K}\hat{R}_{i}^{2}<\sum_{i=1}^{K}(R_{i}^{2})^{\star}$, indicating that there is at least one type $\theta_{i}^{2}$ satisfying $\hat{R}_{i}^{2}<(R_{i}^{2})^{\star},\>i\in\mathcal{K}$.
@@
$\hat{R}_{i}^{2}\geq\hat{R}_{i-1}^{2}+\frac{c(T_{i}^{2}(\theta_{k}^{1})-T_{i-1}^{2}(\theta_{k}^{1}))}{\theta_{i}^{2}}.$ (A-1)
@@
$(R_{i}^{2})^{\star}-(R_{i-1}^{2})^{\star}=\frac{c(T_{i}^{2}(\theta_{k}^{1})-T_{i-1}^{2}(\theta_{k}^{1}))}{\theta_{i}^{2}}.$ (A-2)
By substituting (A-2) into (A-1), we have
$\hat{R}_{i}^{2}-(R_{i}^{2})^{\star}\geq\hat{R}_{i-1}^{2}-(R_{i-1}^{2})^{\star}.$ (A-3)
@@
Appendix B Proof for Lemma 7

### v2

@@ VI-E Performance Evaluation of the EDMSAC Algorithm
Figure 8: Performance of the EDMSAC algorithm in optimal dynamic contract design under critical hyperparameter settings, where the optimal value corresponding to the optimal hyperparameter is marked in red.
As shown in Fig. 8, we present the performance of the EDMSAC algorithm in optimal dynamic contract design with different hyperparameter configurations. Figure 8(a) illustrates the performance of the EDMSAC algorithm under different pruning rates, i.e., $30\%$, $50\%$, and $70\%$. We observe that the EDMSAC algorithm demonstrates superior performance under higher pruning rates, leading to faster convergence and higher test rewards. The reason is that a higher pruning rate indicates that more redundant and unimportant neurons in DM-based actor networks are structurally masked. As a result, the pruned DM-based actor networks exhibit better generalization capabilities and are more effective at learning essential features in optimal dynamic contract design. Figure 8(b) presents the performance of the EDMSAC algorithm under different denoising steps, i.e., $3$, 66, and $12$. We observe that the EDMSAC algorithm achieves faster convergence and yields higher total profits for the cloud server under a smaller denoising step. The reason is that a smaller denoising step can achieve higher computational efficiency of DM-based actor networks. Figure 8(c) illustrates the performance of the EDMSAC algorithm under two groups of different learning rates. The first group is the actor networks with a learning rate of $2\times 10^{-7}$ and the critic networks with a learning rate of $2\times 10^{-6}$. The second group is the actor networks and critic networks with the same learning rate of $10^{-4}$. We observe that the EDMSAC algorithm has better performance under the first group of learning rates, indicating that the pruned actor networks need to learn an efficient policy in optimal dynamic contract design under small learning rates.
VI-F Performance Evaluation of the Dynamic Contract Scheme
In Fig. 9, we illustrate the two-period dynamic contracts generated by the EDMSAC algorithm under three network states, with a pruning rate of $0.5$ and a denoising step of $6$. Additionally, the learning rates of the actor and critic networks are set to $2\times 10^{-7}$ and $2\times 10^{-6}$, respectively. Attributed to the exploration and generalization capabilities of GDMs, the proposed EDMSAC algorithm demonstrates the capability to generate optimal dynamic contracts across varying network states, indicating that our algorithm possesses strong adaptability and generalization abilities to the dynamic environments inherent in federated AI agent construction, thereby ensuring robust system performance in mobile metaverses. We observe that in period 2, the ESs with higher types perform more training rounds $T_{2}^{k}$ for agent module creation and consequently receive higher rewards $R_{2}^{k}$, which validates Lemmas 1 and 2, respectively. In period 1, the ESs with higher types also perform more training rounds $T_{i}^{k}$ for agent module creation and receive higher rewards $R_{i}^{k}$, which demonstrates Lemmas 7 and 8, respectively.
@@ VI-F Performance Evaluation of the Dynamic Contract Scheme
In Fig. 10, we compare the proposed two-period dynamic contract scheme with three baseline contract schemes. Figure 10(a) presents the test reward comparison between the proposed scheme and three baseline contract schemes. Our dynamic contract scheme consistently achieves higher test rewards compared with the other three contract schemes, indicating that the proposed two-period dynamic contract scheme possesses superior performance. In addition, the test rewards obtained by dynamic contract schemes consistently surpass those of static contract schemes, highlighting the advantages of incorporating adaptability in contract design. The reason is that, unlike static contract schemes, dynamic contract schemes take into account the profit of the cloud server in the second period. As a result, the total profit of the cloud server achieved under the dynamic contract scheme is higher than that of the static contract scheme. Figure 10(b) illustrates the relationship between the total profit of the cloud server and varying numbers of ESs under four contract schemes. No matter which contract scheme, as the number of ESs increases, the total profit of the cloud server also increases. Moreover, compared with the three baseline contract schemes, the proposed two-period dynamic contract scheme can achieve the highest total profit for the cloud server. This is primarily because the proposed scheme is capable of designing suitable contracts for ESs based on their respective types rather than generating contracts randomly, which enables more effective resource utilization for federated AI agent construction and contributes to higher total profits for the cloud server. Figure 10(c) presents the impacts of the scaling factor $\alpha$ and the discount factor $\beta$ on the total profit of the cloud server under different numbers of ESs. When $\beta$ remains constant, the total profit of the cloud server increases with the rise of $\alpha$. The reason is that a larger $\alpha$ reflects a greater overall magnitude of economic benefits allocated to the cloud server, thereby enhancing its total profit. In addition, when $\alpha$ remains constant, the total profit of the cloud server increases as $\beta$ increases. This is primarily because the cloud server obtains higher profits in the second period of federated AI agent construction, thus enhancing its total profit.
VII Conclusion and Future Work
@@
In this paper, we have studied federated AI agent construction in mobile metaverses. Specifically, we have proposed an edge-cloud collaboration-based federated AI agent construction framework in mobile metaverses. In this framework, ESs serve as agent infrastructures to construct agent modules in a distributed manner, and the cloud server subsequently integrates them into complete AI agents and deploys the constructed AI agents at the edge, thereby providing AI agent services for users in mobile metaverses. Considering the dynamic information asymmetry in federated AI agent construction, we have designed a two-period dynamic contract model to incentivize ESs to continuously participate in federated AI agent construction. Furthermore, to effectively identify optimal dynamic contracts, we have proposed the EDMSAC algorithm, in which the redundant neurons of DM-based actor networks are dynamically pruned, thereby enhancing denoising efficiency for dynamic contract sampling. Extensive simulations have demonstrated that the proposed EDMSAC can achieve up to a $23\%$ improvement over the DMSAC algorithm in optimal dynamic contract generation. For future work, we aim to develop a mobile AI agent paradigm built upon the proposed framework, which will empower AI agents with enhanced adaptability and decision-making capabilities in real-world mobile environments. Furthermore, we plan to build a working prototype using real edge devices to transform the proposed framework into a deployable system.
Appendix A Proof for Lemma 1
$\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{j}^{2}(\theta_{k}^{1})-cT_{j}^{2}(\theta_{k}^{1}).$ (A-1)
We transform the above inequality into
$\theta_{i}^{2}[R_{i}^{2}(\theta_{k}^{1})-R_{j}^{2}(\theta_{k}^{1})]\geq c[T_{i}^{2}(\theta_{k}^{1})-T_{j}^{2}(\theta_{k}^{1})].$ (A-2)
Since $T_{i}^{2}(\theta_{k}^{1})\geq T_{j}^{2}(\theta_{k}^{1})$, we have $\theta_{i}^{2}[R_{i}^{2}(\theta_{k}^{1})-R_{j}^{2}(\theta_{k}^{1})]\geq 0$, then $R_{i}^{2}(\theta_{k}^{1})\geq R_{j}^{2}(\theta_{k}^{1})$. Similarly, given that $R_{i}^{2}(\theta_{k}^{1})\geq R_{j}^{2}(\theta_{k}^{1})$, we have $T_{i}^{2}(\theta_{k}^{1})\geq T_{j}^{2}(\theta_{k}^{1})$. Thus, the proof of Lemma 1 is completed.
Appendix B Proof for Lemma 2
Based on the IC constraints in (14), we have
$\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{j}^{2}(\theta_{k}^{1})-cT_{j}^{2}(\theta_{k}^{1}),$ (B-1)
$\theta_{j}^{2}R_{j}^{2}(\theta_{k}^{1})-cT_{j}^{2}(\theta_{k}^{1})\geq\theta_{j}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1}).$ (B-2)
By combining (B-1) and (B-2), we have
$(\theta_{i}^{2}-\theta_{j}^{2})(R_{i}^{2}(\theta_{k}^{1})-R_{j}^{2}(\theta_{k}^{1}))\geq 0.$ (B-3)
Since $\theta_{i}^{2}(\theta_{k}^{1})\geq\theta_{j}^{2}(\theta_{k}^{1})$, we have $R_{i}^{2}(\theta_{k}^{1})\geq R_{j}^{2}(\theta_{k}^{1})$. Thus, the proof of Lemma 2 is completed.
Appendix C Proof for Lemma 4
$\begin{split}\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})&-E\geq\theta_{i}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E\\ &\geq\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E\geq 0.\end{split}$ (C-1)
From (C-1), we can observe that if the type-$1$ ES with the lowest WTP has a non-negative utility, it implies that the other IR constraints automatically hold. Additionally, for $\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E\geq 0$, the cloud server would reduce the reward $R_{1}^{2}(\theta_{k}^{1})$ as much as possible until $\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E=0$  [31], thus maximizing its utility. Thus, the proof of Lemma 4 is completed.
Appendix D Proof for Lemma 5
$\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1}).$ (D-1)
$\theta_{i-1}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\theta_{i-1}^{2}R_{i-2}^{2}(\theta_{k}^{1})-cT_{i-2}^{2}(\theta_{k}^{1}),$ (D-2)
where (D-2) can be transformed into
$\theta_{i-1}^{2}(R_{i-1}^{2}(\theta_{k}^{1})-R_{i-2}^{2}(\theta_{k}^{1}))\geq c(T_{i-1}^{2}(\theta_{k}^{1})-T_{i-2}^{2}(\theta_{k}^{1})).$ (D-3)
Based on the ascending order of types, we can further obtain
$\theta_{i}^{2}(R_{i-1}^{2}(\theta_{k}^{1})-R_{i-2}^{2}(\theta_{k}^{1}))\geq c(T_{i-1}^{2}(\theta_{k}^{1})-T_{i-2}^{2}(\theta_{k}^{1})),$ (D-4)
which can be transformed into
$\theta_{i}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{i-2}^{2}(\theta_{k}^{1})-cT_{i-2}^{2}(\theta_{k}^{1}).$ (D-5)
$\begin{split}\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq&\theta_{i}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\\ &\cdots\geq\theta_{i}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1}).\end{split}$ (D-6)
$\begin{split}\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq&\theta_{i}^{2}R_{i+1}^{2}(\theta_{k}^{1})-cT_{i+1}^{2}(\theta_{k}^{1})\geq\\ &\cdots\geq\theta_{i}^{2}R_{K}^{2}(\theta_{k}^{1})-cT_{K}^{2}(\theta_{k}^{1}).\end{split}$ (D-7)
Thus, the proof of Lemma 5 is completed.
Appendix E Proof for Lemma 6
$\theta_{i}^{2}(R_{i}^{2}(\theta_{k}^{1})-R_{i-1}^{2}(\theta_{k}^{1}))\geq\theta_{i-1}^{2}(R_{i}^{2}(\theta_{k}^{1})-R_{i-1}^{2}(\theta_{k}^{1})),$ (E-1)
which (E-1) can be further transformed into
$c(T_{i}^{2}(\theta_{k}^{1})-T_{i-1}^{2}(\theta_{k}^{1}))\geq\theta_{i-1}^{2}(R_{i}^{2}(\theta_{k}^{1})-R_{i-1}^{2}(\theta_{k}^{1})),$ (E-2)
and we have
$\theta_{i-1}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\theta_{i-1}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1}).$ (E-3)
Therefore, the LUIC constraints remain valid. The proof of Lemma 6 is completed.
Appendix F Proof for Theorem 2
@@
We use the contradiction method to demonstrate the uniqueness of the optimal reward in period 2. We assume that there is another optimal reward $\hat{R}_{i}^{2}$ that facilitates the greater profit of the cloud server. Since the utility of the cloud server is inversely proportional to the total reward, we can obtain $\sum_{i=1}^{K}\hat{R}_{i}^{2}<\sum_{i=1}^{K}(R_{i}^{2})^{\star}$, indicating that there is at least one type $\theta_{i}^{2}$ satisfying $\hat{R}_{i}^{2}<(R_{i}^{2})^{\star},\>i\in\mathcal{K}$ [14, 7].
@@
$\hat{R}_{i}^{2}\geq\hat{R}_{i-1}^{2}+\frac{c(T_{i}^{2}(\theta_{k}^{1})-T_{i-1}^{2}(\theta_{k}^{1}))}{\theta_{i}^{2}}.$ (F-1)
@@
$(R_{i}^{2})^{\star}-(R_{i-1}^{2})^{\star}=\frac{c(T_{i}^{2}(\theta_{k}^{1})-T_{i-1}^{2}(\theta_{k}^{1}))}{\theta_{i}^{2}}.$ (F-2)
By substituting (F-2) into (F-1), we have
$\hat{R}_{i}^{2}-(R_{i}^{2})^{\star}\geq\hat{R}_{i-1}^{2}-(R_{i-1}^{2})^{\star}.$ (F-3)
@@
Appendix G Proof for Lemma 7